OpenStudy (anonymous):

the sum of three numbers is 19. the sum of twice the first number, 3 times the second number, and four times the third number is 61. the difference between 4 times the first number and second number is 9. Find all three numbers

OpenStudy (anonymous):

x+y+z=19 2x+3y+4z=61 4x-y=9

OpenStudy (anonymous):

i got x=-12 y=-57 z=88

OpenStudy (anonymous):

i got x=4 , y=7, z=8

OpenStudy (anonymous):

how

OpenStudy (anonymous):

x+y+z=19 ------------multiply by -4 2x+3y+4z=61

OpenStudy (anonymous):

ok i did that

OpenStudy (anonymous):

-2x-y=-15

OpenStudy (anonymous):

-2x-y=-15 4x-y=9 ---------------subtract

OpenStudy (anonymous):

or multiply by -1 then add

OpenStudy (anonymous):

@Mandy86 got it?

OpenStudy (anonymous):

yes thank you can u also check my work

OpenStudy (anonymous):

on my other question

OpenStudy (anonymous):

i did, good job